d) The unit–vector notation of the two vectors a = 4.0i + 3.0j and b = −13.0i + 7.0j. Determine




the magnitude and direction of a + b [4mks

1
Expert's answer
2021-12-14T21:32:28-0500

Sum of the vectors:

"a + b = (4.0 + (-13.0))i + (3.0 + 7.0)j = -9.0i + 10.0j"



Magnitude:

"|| a + b || = \sqrt{(-9.0)^2 + 10.0^2} \approx 13.45"



Direction:

"\tan(\theta) = \cfrac{10.0}{-9.0} \\
\theta = \tan^{-1} (\cfrac{10.0}{-9.0}) \approx 132.0 \degree"